OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

OpenStudy (hoblos):

general formula of point-slope form is y-b=m(x-a) where m is the slope and (a,b) is a point on the line

OpenStudy (hoblos):

so y+1=6(x-5)

OpenStudy (anonymous):

thanks(:

OpenStudy (hoblos):

my pleasure :)